# Generative-Enhanced Heterogeneous Graph Contrastive Learning

### v2

@@ Abstract.
Heterogeneous Graphs (HGs) can effectively model complex relationships in the real world by multi-type nodes and edges. In recent years, inspired by self-supervised learning, contrastive Heterogeneous Graphs Neural Networks (HGNNs) have shown great potential by utilizing data augmentation and contrastive discriminators for downstream tasks. However, data augmentation is still limited due to the graph data’s integrity. Furthermore, the contrastive discriminators remain sampling bias and lack local heterogeneous information. To tackle the above limitations, we propose a novel Generative-Enhanced Heterogeneous Graph Contrastive Learning (GHGCL). Specifically, we first propose a heterogeneous graph generative learning enhanced contrastive paradigm. This paradigm includes: 1) A contrastive view augmentation strategy by using a masked autoencoder. 2) Position-aware and semantics-aware positive sample sampling strategy for generating hard negative samples. 3) A hierarchical contrastive learning strategy for capturing local and global information. Furthermore, the hierarchical contrastive learning and sampling strategies aim to constitute an enhanced contrastive discriminator under the generative-contrastive perspective. Finally, we compare our model with seventeen baselines on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on node classification and link prediction tasks.
@@
Self-Supervised Learning, Heterogeneous Graph Neural Networks, Contrastive Learning, Generative Learning
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligence; Neural networks
1. INTRODUCTION
@@
Heterogeneous Information Networks (HIN) or Heterogeneous Graphs (HG) (Sun and Han, 2013; Wang et al., 2019, 2021; Shi et al., 2018; Tian et al., 2023) have become critical tools for analyzing the real world, as they preserve rich semantic information by modeling different types of nodes and edges. This is particularly relevant in networks such as citation networks (Dong et al., 2017; Wang et al., 2019), recommendation systems (Shi et al., 2018; Chen et al., 2023), and social networks (Long et al., 2021). Recently, Heterogeneous Graph Neural Networks (HGNNs) (Wang et al., 2019, 2021; Hu et al., 2020a; Tian et al., 2023) have demonstrated a strong ability in processing HIN data because they have expanded the receptive field of nodes through meta-path (Sun et al., 2011) and message-passing mechanism, thereby capturing the rich semantics.
Figure 1. (a)Heterogeneous graph self-supervised learning with contrastive paradigms; (b)Heterogeneous graph self-supervised learning with generative-contrastive paradigms.
@@
Self-supervised learning (SSL) has been widely applied in the fields of computer vision (CV) (He et al., 2020b; Chen et al., 2020; Grill et al., 2020) and natural language processing (NLP) (Chowdhery et al., 2023; Kenton and Toutanova, 2019), leveraging unlabeled data to learn informative representations. Inspired by this trend, researchers explore the combination of HGNNs and self-supervised learning. As a major paradigm, contrastive HGNNs (Wang et al., 2021; Zhu et al., 2022; Wang et al., 2023; Chen et al., 2023) based SSL have shown great potential because they are more suitable for downstream tasks by setting contrastive view and discriminator. Data augmentation (Wu et al., 2021; Yang et al., 2023) is typically used as a tool to construct contrastive views, as it maximizes mutual information by aligning two views. However, due to the need to ensure graph data’s integrity and unique features, it is difficult to provide reasonable data augmentation to leverage the advantages of contrastive learning (He et al., 2020b; Liu et al., 2021).
@@
Generative learning (Chowdhery et al., 2023; Tian et al., 2023; Hou et al., 2022; Kenton and Toutanova, 2019; Liu et al., 2021) can effectively solve the limitation because it can reconstruct the data distribution without considering downstream tasks and without data augmentation, allowing each node embedding to be enhanced. This idea also provides a foundation for us to propose a generative learning enhanced contrastive paradigm for heterogeneous graphs. As shown in Figure 1, first, HGNNs utilize a meta-path to obtain a graph containing nodes of the same type. Contrastive learning (Wang et al., 2023; Zhu et al., 2022) generates supervised signals by aligning two similar views constructed through random edge dropping; while generative-contrastive learning (Goodfellow et al., 2020; Yang et al., 2023; Wu et al., 2023) reconstructs the intrinsic structures and features of data by training an autoencoder and uses a contrastive discriminator to distinguish them in the embedding space. Despite the obvious advantages of the generative-contrastive paradigm, there are still some challenges for heterogeneous graphs:
@@
CH1: Unbalanced data augmentation. The current augmentation strategies (Wang et al., 2023; Zhu et al., 2022; Lin et al., 2022; Wu et al., 2021), such as structural and feature augmentation, still have limitations. Structural augmentation (Wang et al., 2023; Zhu et al., 2022; Wu et al., 2021), which involves randomly dropout nodes and edges, could disrupt the graph integrity since nodes and edges are the fundamental properties of a graph. On the other hand, feature augmentation (Wang et al., 2021; Lin et al., 2022), which perturbs node attributes, ignores the unique features of each node’s input. In such methods, some nodes have more edges and embeddings than others, resulting in unbalanced supervised signals. Exploring how to better utilize meta-paths to design enhanced contrastive views remains a challenge that needs to be addressed.
@@
CH2: Negative sampling bias. In heterogeneous graphs, contrastive learning (Wang et al., 2023, 2021; Zhu et al., 2022) heavy relies on high-quality negative samples. Moreover, the contrastive discriminator has stricter requirements for sampling under the generative-contrastive paradigm (Wu et al., 2023). Existing sampling strategies (Wang et al., 2023, 2021; Zhu et al., 2022) are usually based on randomness or meta-paths, which often increase false negative samples. Some recent works (Kalantidis et al., 2020; Liu et al., 2023) aim to select or synthesize hard negatives to mitigate the impact of false negatives. However, they are not suitable for heterogeneous graphs and offer limited benefits. Therefore, designing an enhanced positive sampling strategy for heterogeneous graphs is urgently needed.
@@
CH3: Lack local heterogeneous information. Existing methods (Wang et al., 2023; Zhu et al., 2022; Tian et al., 2023) rely solely on meta-paths to capture rich semantic information in heterogeneous graphs and often overlook the influence of one-hop neighbors on target nodes. Network schema leverages direct connections between nodes to capture the local structure of target nodes. In contrast, meta-paths extract higher-order structures. Designing an effective self-supervised mechanism that considers both local and global information can better help the model work.
To tackle the three challenges mentioned above, we propose a heterogeneous graph self-supervised model called Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). For CH1, GC-HGNN employs a generative masked autoencoder (MAE) to enhance contrastive views. The MAE can reconstruct node embeddings without altering the original graph structure and features. The masking strategy averages the edges of each node to ensure that each node receives the same supervised signals during the reconstruction process. For CH2, the model proposes a sampling strategy that combines location-awareness and semantic-awareness. This approach adaptively samples nodes based on the rich semantics unique to heterogeneous graphs and the global positions of nodes in the heterogeneous graph. Finally, for CH3, GC-HGNN utilizes hierarchical contrastive (HCL) learning to capture one-hop and higher-order neighbors. Intra-contrast is used to contrast different meta-path views, while inter-contrast is used to contrast between network schema and meta-path views. We summarize the contributions of this paper as follows:
@@
  • We propose a heterogeneous graph generative learning enhanced contrastive paradigm, leveraging a generative masked autoencoder as a tool for contrastive view augmentation. It is a meaningful exploration for graph generative-contrastive networks.
@@
  • We introduce hierarchical contrastive learning as an enhanced contrastive discriminator, combined with novel sampling strategies to align with the generative model. These modules can better mine self-supervised signals under an end-to-end training scenario.
@@
  • We conduct comprehensive experiments to evaluate GC-HGNN with seventeen baseline methods and eight real datasets. The experimental results show that GC-HGNN is superior to other methods and effectively extracts heterogeneous information.
2. RELATED WORK
@@
This section discusses the development process of heterogeneous graph neural networks. Here, we introduce the issue of infeasible and expensive data labels, further elaborating on the self-supervised advantages and providing a brief overview of existing models.
2.1. Heterogeneous Graph Neural Networks
@@
In recent years, Graph Neural Networks (GNNs) (You et al., 2020; Hu et al., 2020b; Hou et al., 2022; Veličković et al., 2018; He et al., 2020a) have achieved significant success in fields such as social network analysis and recommender systems. This success relies on their effective modeling of graph structure information and local and global information integration. These graph structures often have different types of nodes and edges. Heterogeneous graph neural networks are designed to better handle heterogeneous graph data.
@@
The currently popular heterogeneous graph neural network models (Wang et al., 2019, 2021; Zhang et al., 2019; Fu et al., 2020; Yu et al., 2022b) can be classified into meta-path-based methods and types-based methods. The classic meta-path-based models include HAN (Wang et al., 2019), HetGNN (Zhang et al., 2019), and MAGNN (Fu et al., 2020). Specifically, HAN employs a dual-layer attention mechanism at both the node and semantic levels to adaptively capture the relationships between nodes and meta-paths. HetGNN introduces a heterogeneous graph attention propagation mechanism, combining meta-path information and attention mechanisms to learn node representations. MAGNN integrates meta-paths and adaptive mechanisms to learn node representations in heterogeneous graphs by learning meta-path attention weights at both the node and global levels. The type-based models include RGCN (Schlichtkrull et al., 2018), HGT (Hu et al., 2020a), and RHGNN (Yu et al., 2022b). RGCN captures rich semantic relationships by using different weight matrices for different types of edges. HGT achieves efficient representation learning and information propagation on heterogeneous graph data by introducing a multi-head attention mechanism and a heterogeneous graph structure encoder. RHGNN utilizes dedicated graph convolution components and cross-relation passing modules to learn node representations on heterogeneous graphs. However, the methods mentioned above only investigate supervised heterogeneous graph models. In order to deeply explore the potential relationships and patterns in the data and to effectively utilize unlabeled data, an increasing number of studies are beginning to focus on self-supervised learning methods.
2.2. Self-supervised Graph Learning
@@
With the rapid development of self-supervised learning in computer vision (He et al., 2020b; Chen et al., 2020) and natural language processing (Kenton and Toutanova, 2019; Chowdhery et al., 2023) fields, graph neural networks have drawn inspiration from this progress. Self-supervised graph neural networks often train without labels by generating contrastive views or reconstructing node attributes, thus better leveraging the rich structural information in graph data. Transitioning from general graphs to heterogeneous graphs involves assigning different types or attributes to nodes and edges in the graph, making the graph structure more complex and diverse in representation.
@@
Self-supervised heterogeneous graph neural networks can be categorized into contrastive and generative. Contrastive learning models (Ren and Liu, 2020; Wang et al., 2021, 2023) generate enhanced contrastive views and attempt to maximize the mutual information between these views to produce supervised signals. Specifically, HDGI (Ren and Liu, 2020) utilizes graph convolutional modules and semantic-level attention to learn node representations and maximize local-global mutual information. HeCo (Wang et al., 2021) complements across perspectives through network schema and meta-path views, establishing a sampling contrast method between the two views to achieve self-supervised enhancement. HGCML (Wang et al., 2023) solely utilizes the rich semantic relationships brought by meta-paths for graph enhancement, introducing multi-view contrastive learning. It also proposes sampling strategies for positive samples in both topology and semantics to alleviate sampling bias. Generative models (Zheng et al., 2021; Tian et al., 2023) commonly adopt graph autoencoder and attempt to reconstruct features to obtain supervision. For example, MvDGAE (Zheng et al., 2021) utilizes random edge dropout to enhance views and autoencoders to enforce the recovery of views for more robust representations. Meanwhile, HGMAE (Tian et al., 2023) further introduces a dynamic masking mechanism and position-aware encoding to enhance the model’s performance and utilizes attribute and view dual-recovery to improve denoising capability.
Figure 2. A toy example of an HIN on ACM (Wang et al., 2019) dataset.
3. PRELIMINARIES

### v3

# Generative-Contrastive Heterogeneous Graph Neural Network

@@ Abstract
Heterogeneous Graphs (HGs) effectively model complex relationships in the real world through multi-type nodes and edges. In recent years, inspired by self-supervised learning (SSL), contrastive learning (CL)-based Heterogeneous Graphs Neural Networks (HGNNs) have shown great potential in utilizing data augmentation and contrastive discriminators for downstream tasks. However, data augmentation remains limited due to the graph data’s integrity. Furthermore, the contrastive discriminators suffer from sampling bias and lack local heterogeneous information. To tackle the above limitations, we propose a novel Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). Specifically, we propose a heterogeneous graph generative learning method that enhances CL-based paradigm. This paradigm includes: 1) A contrastive view augmentation strategy using a masked autoencoder. 2) Position-aware and semantics-aware positive sample sampling strategy for generating hard negative samples. 3) A hierarchical contrastive learning strategy aimed at capturing local and global information. Furthermore, the hierarchical contrastive learning and sampling strategies aim to constitute an enhanced contrastive discriminator under the generative-contrastive perspective. Finally, we compare our model with seventeen baselines on eight real-world datasets. Our model outperforms the latest baselines on node classification and link prediction tasks. To reproduce our work, we have open-sourced our code at https://github.com/wangyu0627/GC-HGNN.
Index Terms:
@@
Self-Supervised Learning, Heterogeneous Graph Neural Networks, Contrastive Learning, Generative Learning.
I Introduction
@@
Heterogeneous Information Networks (HIN) or Heterogeneous Graphs (HG) [1, 2] have emerged as critical tools for analyzing the real world, as they capture rich semantic information by modeling different types of nodes and edges. This is particularly relevant in networks such as citation networks [3, 2], recommender systems [4, 5], and social networks [6]. Recently, Heterogeneous Graph Neural Networks (HGNNs) [2, 7] demonstrate a strong ability in processing HIN data because they expand the receptive field of nodes through meta-path [8] and message-passing mechanism, thereby capturing the rich semantics.
Figure 1: (a) Contrastive Learning paradigm aligns two similar views constructed through random edge dropping; (b) Generative-Contrastive paradigm aligns two reconstructed embeddings without altering the original graph structure.
@@
Self-supervised learning (SSL) has been widely applied in the fields of Computer Vision (CV) [9, 10] and Natural Language Processing (NLP) [11, 12] leveraging unlabeled data to learn informative representations. Inspired by this trend, researchers explore the combination of HGNNs and SSL. As a major paradigm, HGNNs [7, 13] based on Contrastive Learning (CL) show great potential, primarily due to their suitability for downstream tasks achieved by setting contrastive views and discriminators, as demonstrated in [9, 14]. Data augmentation [15, 16] serves as a tool to construct construct contrastive views, maximizing mutual information by aligning two views. However, given the requirement to maintain graph data’s integrity and unique features, it is challenging to provide reasonable data augmentation to leverage the benefits of contrastive learning.
@@
Generative learning [11, 12] effectively addresses this limitation by reconstructing the data distribution, disregarding downstream tasks and data augmentation, thereby enhancing each node embedding. This idea further serves as a foundation for proposing a generative learning enhanced contrastive learning paradigm for heterogeneous graphs. As shown in Figure 1, first, HGNNs utilize a meta-path to obtain a graph containing nodes of the same type. CL-based HGNN [17, 13] generates supervised signals by aligning two similar views constructed through random edge dropping; meanwhile, generative-contrastive learning [18, 19] reconstructs the intrinsic structures and features of data by training an autoencoder and uses a contrastive discriminator to distinguish them in the embedding space. Despite the obvious advantages of the generative-contrastive paradigm, it presents several challenges for heterogeneous graphs:
@@
CH1: Unbalanced data augmentation. The current graph augmentation strategies [17, 15], such as structural and feature augmentation, still have limitations. On the one hand, structural augmentation [17, 13], which involves randomly dropping out nodes and edges, might disrupt the graph integrity, as nodes and edges are the fundamental properties of a graph. On the other hand, feature augmentation [7, 20], which perturbs node attributes, tends to overlook the unique features of each node’s input. In these methods, certain nodes have more edges than others, which results in unbalanced supervised signals. Exploring how to design enhanced contrastive views represents a challenge that requires attention.
@@
CH2: Negative sampling bias. In heterogeneous graphs, contrastive learning [7] heavily relies on high-quality negative samples. Moreover, the contrastive discriminator requires stricter sampling methods under the generative-contrastive paradigm [19]. Existing sampling strategies [17, 7] are typically based on randomness or meta-paths, and these approaches frequently lead to an increase in false negative samples. Recent works [21, 22] aim to select or synthesize hard negatives to mitigate the impact of false negatives. However, they are not suitable for HGs and provide limited benefits. Therefore, there is an urgent need to design an enhanced sampling strategy for HGs.
@@
CH3: Lack local heterogeneous information. Existing methods [17, 23] rely solely on meta-paths to capture rich semantic information in HGs and often overlook the influence of one-hop neighbors on target nodes. Network schema leverages direct connections between nodes to capture the local structure of target nodes. In contrast, meta-paths are employed to extract higher-order structures. Designing an effective self-supervised mechanism that considers both local and global information can significantly enhance the model’s performance.
To address the three challenges mentioned above, we propose a heterogeneous graph self-supervised model called Generative-Contrastive Heterogeneous Graph Neural Network (GC-HGNN). For CH1, GC-HGNN employs a generative masked autoencoder (MAE) to enhance contrastive views. The MAE is capable of reconstructing node embeddings without altering the original graph structure and features. The masking strategy averages the edges of each node, ensuring that each node receives the same supervised signals during the reconstruction process. Moreover, random structural perturbations in existing methods can exacerbate imbalance, potentially leading to adverse effects on node representations [16]. Our proposed GC-HGNN uses an autoencoder to create embeddings without modifying the graph’s topology or inherent node features, preserving useful information. The proposed sampling strategy is not simply ”node drop” but a selective uniform random sampling without replacement. This method helps prevent bias by ensuring neighbors are neither fully masked nor entirely visible, which enhances feature restoration [24]. For CH2, the model introduces a sampling strategy that integrates both location-awareness and semantic-awareness. This approach adaptively samples nodes based on the rich semantics inherent to HGs and the global positions of nodes within these graphs. In contrastive learning, the goal is to minimize positive sample distances while maximizing negative ones. In GC-HGNN, the positive sampling strategy also improves hard negative selection, promoting feature similarity in positive pairs while reducing similarity in negative pairs. Studies show that suboptimal false negatives can lead to erroneous guidance [25], and precise positive sampling enhances discrimination, thus indirectly supporting negative sample quality [26]. Finally, for CH3, GC-HGNN employs hierarchical contrastive learning (HCL) to capture both one-hop and higher-order neighbors. Intra-contrast is employed to contrast different meta-path views, while inter-contrast is used to contrast between network schema and meta-path views. Our model builds upon HeCo’s framework but simplifies and generalizes it by encoding all edge types directly, which minimizes preprocessing cost and avoids semantic loss. Unlike HeCo’s two-step attention aggregation, our approach is more streamlined.
We summarize the contributions of this paper as follows:
@@
  • We propose a generative learning method that enhances contrastive learning (CL)-based paradigm GC-HGNN, leveraging a generative masked autoencoder for view augmentation. This represents a significant exploration for graph generative-contrastive networks.
@@
  • We introduce hierarchical contrastive learning (HCL) as an enhanced contrastive discriminator, which is integrated with innovative sampling strategy to align the generative model. These modules effectively mine self-supervised signals.
@@
  • We conduct comprehensive experiments to evaluate GC-HGNN with seventeen baseline methods and eight real datasets. The experimental results demonstrate that GC-HGNN is superior to baselines and effectively extracts heterogeneous information.
Related Work
@@
This section examines the development process of heterogeneous graph neural networks. Here, we introduce the issue of infeasible and expensive data labels, further expanding on the self-supervised advantages and providing a brief overview of existing models.
I-A Heterogeneous Graph Neural Networks
@@
In recent years, Graph Neural Networks (GNNs) [27, 28, 29, 30] have demonstrated significant success in fields such as social network analysis and recommender systems. This success is attributed to their effective modeling of graph structure information and the integration of both local and global information. These graph structures typically have different types of nodes and edges. Heterogeneous graph neural networks (HGNNs) are designed to better handle heterogeneous information.
@@
The currently prevalent HGNN-based models [31, 7, 32, 33, 34] can be categorized into meta-path-based methods and type-based methods. The classic meta-path-based models include HAN [2], HetGNN [32], and MAGNN [33]. Specifically, HAN employs a dual-layer attention mechanism at both the node and semantic levels to adaptively capture the relationships between nodes and meta-paths. HetGNN introduces a heterogeneous graph attention propagation mechanism, combining meta-path information and attention mechanisms to learn node representations. MAGNN integrates meta-paths and adaptive mechanisms to learn node representations in heterogeneous graphs by learning meta-path attention weights at both the node and global levels. The type-based models include RGCN [35], HGT [36], and RHGNN [34]. RGCN captures rich semantic relationships by using different weight matrices for different types of edges. HGT achieves efficient representation learning and information propagation on heterogeneous graph data by introducing a multi-head attention mechanism and a heterogeneous graph structure encoder. RHGNN utilizes dedicated graph convolution components and cross-relation passing modules to learn node representations on heterogeneous graphs. However, the methods mentioned above focus exclusively on supervised heterogeneous graph models. FrameERC [37, 38] leverages the Framelet Transform in a multimodal GNN for emotion recognition in conversations. In order to deeply explore the potential relationships and patterns in the data and to effectively utilize unlabeled data, an increasing number of studies are beginning to focus on self-supervised learning methods.
I-B Self-supervised Graph Learning
@@
Given the rapid development of self-supervised learning in computer vision [9, 39] and natural language processing [12, 11] fields, graph neural networks have benefited from this progress. Self-supervised graph neural networks typically train without labels by generating contrastive views or reconstructing node attributes, thereby better leveraging the rich structural information in graph data. Transitioning from homogeneous graphs to heterogeneous graphs entails assigning different types or attributes to nodes and edges in the graph, rendering the graph structure more complex and diverse in representation.
@@
Self-supervised heterogeneous graph neural networks can be categorized into contrastive and generative. Contrastive learning models [40, 7, 17] generate enhanced contrastive views and aim to maximize the mutual information between these views to produce supervised signals. Specifically, HDGI [40] utilizes graph convolutional modules and semantic-level attention to learn node representations and maximize local-global mutual information. HeCo [7] complements across perspectives through network schema and meta-path views, establishing a sampling contrast method between the two views to achieve self-supervised enhancement. HGCML [17] solely utilizes the rich semantic relationships brought by meta-paths for graph enhancement, introducing multi-view contrastive learning. It also proposes sampling strategies for positive samples in both topology and semantics to alleviate sampling bias. Generative models [41, 23] commonly adopt graph autoencoder and aim to reconstruct features to obtain supervision. For example, MvDGAE [42] utilizes random edge dropout to enhance views and autoencoders to enforce the recovery of views for more robust representations. Meanwhile, HGMAE [23] further introduces a dynamic masking mechanism and position-aware encoding to enhance the model’s performance and utilizes attribute and view dual-recovery to improve denoising capability.
Figure 2: A toy example of an HIN on ACM dataset.
TABLE I: Summary of symbols and notations.
Symbol Description
$\mathcal{G}$ A heterogeneous information network (HIN)
$V$, $E$ Sets of nodes and edges
$I$, $J$ Sets of node types and edge types
$T_{G}$ Network schema of the HIN
$\rho_{j}$ A meta-path in the set of selected meta-paths
$\boldsymbol{P}_{r}$ Projection matrix
$\alpha_{ew}$ Attention weight between node $e$ and neighbor $w$
$\sigma$ Activation function
$\boldsymbol{h}_{e}^{1}$ Updated embedding of node $e$ in network schema
$G=\mathcal{V},\mathcal{A},\boldsymbol{X}$ Definition of a meta-path-based-subgraph with
nodes, adjacency matrix, and feature matrix
$\mathcal{V}^{[Mask]}$ Set of masked nodes
$p$ Mask ratio for masked nodes
$\widetilde{\boldsymbol{X}}$ Masked node feature matrix
$\mathcal{V}^{[Remask]}$ Set of remasked nodes
$\boldsymbol{H},\widetilde{\boldsymbol{H}},\boldsymbol{H}_{Re}$ Encoded feature, decoder input with the remasked
feature matrix, reconstructed output
$w_{\rho_{j}}$ Weight score for meta-path $\rho_{j}$
$\gamma_{\rho_{j}}$ Attention weight for meta-path $\rho_{j}$
$\boldsymbol{h}_{e}^{2}$ Updated embedding of node $e$ in meta-path view
$\mathcal{E}$ Edge sampling set for inter-contrast
$\widetilde{\mathcal{V}}$ Positive sample set for intra-contrast
$\mathbb{P}_{i}$, $\mathbb{N}_{i}$ Positive and negative sample sets for inter-contrast
$\tau$ Temperature parameter in contrastive loss
$\lambda$ Balance coefficient of contrastive loss $\mathcal{L}_{Inter}$
$\mathcal{L}_{Gen/Intra/Inter}$ Loss functions of generative learning,
intra-contrast and inter-contrast
$\mathcal{L}_{bpr}$ Bayesian personalized ranking loss
II PRELIMINARIES